Circular No. 8984

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2008 L9–L15 AND C/2008 M1–M7 (SOHO) Further to IAUC 8983, additional near-sun presumed comets have been

found on SOHO website images — all Kreutz sungrazers except for C/2008 L11 and C/2008 L15 (members of the Meyer group). K. Battams writes that C/2008 L9 (which was also found by B. Zhou) was small and stellar in appearance (mag ∼ 7) in C3 images, and quite diffuse with a short, diffuse tail in C2 images. C/2008 L10 and C/2008 L15 were stellar in appearance (mag ∼ 7–7.5). C/2008 L11 was stellar in appearance, reaching mag ∼ 5.5–6 in C2 images. C/2008 L12 was quite diffuse (mag ∼ 7.5–8). C/2008 L13, C/2008 M2, and C/2008 M5 were slightly diffuse (mag ∼ 8). C/2008 L14 was teardrop-shaped (mag ∼ 3–4) in C3 images, and showed a diffuse tail ∼ 0^{o}.25 long in C2 images (the tail also similarly visible in SECCHI COR2-A images, and as somewhat fainter and shorter in COR2-B images). C/2008 M1 was tiny and stellar in appearance (mag ∼ 8). C/2008 M3 was slightly diffuse (mag ∼ 7.5–8). C/2008 M4 was extremely faint (mag ∼ 8.5) and very diffuse. C/2008 M6 was very small, faint (mag ∼ 8), and somewhat stellar in appearance. C/2008 M7 (which was also found by H. Su) was very diffuse (mag ∼ 8).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 L9 June 10.971 5^{h}14.^{m}2 +20^{o}38′ C3/2 MU 2008-O23
C/2008 L10  11.988 5 18.1 +20 38 C3/2 MU 2008-O24
C/2008 L11  12.438 5 19.1 +21 16 C2/3 MK 2008-O24
C/2008 L12  13.314 5 23.1 +21 21 C2  MK 2008-O24
C/2008 L13  13.454 5 23.8 +21 21 C2  MK 2008-O24
C/2008 L14  15.138 5 31.4 +18 59 C3/2* TH 2008-O27
C/2008 L15  15.746 5 35.3 +24 04 C2  RK 2008-O27
C/2008 M1  21.660 5 55.8 +21 39 C2  MK 2008-O27
C/2008 M2  22.463 5 58.7 +21 41 C2  MK 2008-O27
C/2008 M3  23.811 6 03.9 +21 49 C2  HS 2008-O27
C/2008 M4  25.285 6 10.1 +21 40 C2  MK 2008-P33
C/2008 M5  25.935 6 12.3 +21 44 C2  RM 2008-P33
C/2008 M6  27.826 6 20.0 +21 44 C2  TH 2008-P33
C/2008 M7  28.826 6 24.3 +21 45 C2  TH 2008-P33

COMET C/2008 N1 (HOLMES) Corrigendum. On IAUC 8959, lines 2–3, for 0.40-m f/5.8 Schmidt-

Cassegrain reflector read 0.61-m f/4 astrograph

2008 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green